# Corresponding Author

### v3

@@ 3.1.4 Prompt Engineering
The construction of training prompts for samples typically starts with adding prefix words before the modifier token. A simplest example is “Photo of V*”. However, DreamBooth 2ruiz2023dreambooth noted that such a simple description causes a long training time and unsatisfactory performance. To address this, they incorporate the unique modifier with a class noun to describe the SoI in the references (e.g., “Photo of V* cat”). Also, the training caption for each training reference can be more precious for better disentanglement of SoI and irrelevant concepts 57he2023data , such as “Photo of V* cat on the chair”. This follows the tending that high-quality captions in the training set could assist in further improvement of accurate text control BetkerImprovingIG .
@@ 3.2 Pre-trained Adaptation (PTA) Framework
The PTA framework has emerged as a breakthrough approach for PCS, aiming to eliminate the computational burden of per-request fine-tuning while maintaining high-quality, subject-specific generation capabilities. To achieve this goal, this approach combines large-scale pretraining with reference-aware architectures to enable single-pass personalization, as shown in Fig. 5. Based on this architecture, three critical design factors are considered to ensure practical viability: 1) Preservation of semantic-critical features to guarantee visual consistency with reference inputs, 2) Effective fusion that combines reference features with text guidance to achieve desired generation, 3) Optimization of training dataset scale to achieve robust generalization without overfitting. In the following sub-section, we first present overall architecture introduction in Sec. 3.2.1 and then delve into three detailed factors.
3.2.1 Pre-trianed Adaption
@@ 3.2.2 Subject Feature Extraction
Extracting representative features of the SoI is crucial in the creation of personalized content. A common approach is to employ an encoder, leveraging pre-trained models such as CLIP radford2021learning and BLIP li2022blip . While these models excel at capturing global features, they often include irrelevant information that can detract from the fidelity, potentially compromising the quality of the personalized output, such as including the same background in the generation. To mitigate this issue, some studies incorporate additional prior knowledge to guide the learning process so as to focus on the targeted SoI. For instance, the SoI-specific mask 21zhang2023personalize ; 48wang2023hifi ; 74avrahami2023break ; 81safaee2023clic ; 51lu2024object ; 95achlioptas2023stellar contributes to the effective exclusion of the influence of the background. Moreover, using facial landmarks 37wang2024instantid in the context of human face customization helps improve identity preservation. We will discuss a more detailed technical summary of mask-assistant generation in Sec. 4.2.
Handling multiple input references presents another challenge but is essential for real-world deployment. This necessitates an ensemble of features from the multiple reference images to augment the framework’s adaptability. Yet, the majority of current PTA systems are limited to supporting one reference input. Some research works 37wang2024instantid ; 28li2023photomaker propose to average or stack features extracted from multiple references to form a composite SoI representation.
@@ 3.2.3 Subject Feature Fusion
Concatenation-based Fusion 19wei2023elite ; 28li2023photomaker ; 54xiao2023fastcomposer ; 89chen2023dreamidentity ; 85cai2023decoupled ; 96peng2023portraitbooth . This method makes the unique modifier a placeholder token to encapsulate visual subject characteristics. The placeholder token embedding, initialized with image features from visual encoders, is concatenated with text embeddings from the language model. This combined representation subsequently guides generation through standard cross-attention layers in the diffusion process, enabling basic subject-text alignment while maintaining architectural simplicity.
Cross-attention Fusion 40ye2023ip ; 11chen2024subject ; 37wang2024instantid ; 91li2023stylegan ; 212liu2024stylecrafter ; 221zhang2024ssr ; 220personalizationms . This paradigm extends the U-Net architecture with specialized attention mechanisms that jointly process visual and textual conditions. For example, IP-Adapter 40ye2023ip introduces decoupled cross-attention layers that maintain separate query projections for image and text features. In this case, the unique modifier directly displays the plain text.
Multimodal Encoder Fusion 10li2024blip ; 84zhou2023customization ; song2024moma ; 77pan2023kosmos ; 71ma2023unified . This approach leverages powerful multimodal encoder architectures (e.g., BLIP-2 li2023blip ) to jointly embed visual and textual subject descriptors. BLIP-Diffusion 10li2024blip exemplifies this strategy by learning a compact subject prompt embedding that fuses image patches with textual names through Q-Former modules.
Hybrid Fusion 16ma2023subject ; 156zhu2024multibooth . Moreover, some systems integrate multiple fusion strategies. For example, Subject-Diffusion 16ma2023subject integrate both concatenation and cross-attention fusion, taking advantage of the strengths of each approach to enhance the overall personalization capability.
@@ 3.2.4 Training Data
Triplet Data (Reference Image, Target Image, Target Caption). This dataset format is directly aligned with the PCS objectives, establishing a clear relation between the reference and the personalized content. However, such large-scale triplet samples are not widely available. Several strategies have been proposed to mitigate this issue: 1) Data Augmentation. Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data 10li2024blip . 2) Synthetic Sample Generation. Methods like SuTI 11chen2024subject utilize multiple TTF models to generate synthetic samples, which are then paired with original references. 3) Utilizing Recognizable SoIs. Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization 60yuan2023inserting .
Dual Data (Reference Image, Reference Caption). This dataset is essentially a simplified version of the triplet format, where the personalized content is the original image itself. Such datasets are more accessible, including collections like LAION schuhmann2022laion and LAION-FACE zheng2022general . However, a notable drawback is that training tends to focus more on reconstructing the reference image rather than incorporating the text prompts. Consequently, models trained on this type of data might struggle with complex prompts that require substantial modifications or interactions with objects.
@@ 3.3 Hybrid Framework
Recently, some works have started to explore the combination of TTF and PTA methods. HyperDreamBooth 58ruiz2023hyperdreambooth states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details. They first develop a PTA network, followed by a subject-driven fine-tuning. Similarly, DreamTuner 86hua2023dreamtuner pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction. Then an additional fine-tuning stage is conducted for fine identity preservation. In contrast, SuTI 11chen2024subject first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
@@ 4 Techniques in Personalized Content Synthesis
Building upon the architectural framework discussed in Sec. 3, this section analyzes learning optimization techniques applicable to both frameworks. We focus on four categories: attention mechanisms, mask-guided generation, data augmentation, and regularization strategies. These methods aim to address key challenges in PCS, like enhancing subject fidelity, minimizing the interference of redundant semantics, enhancing generalization, and avoiding overfitting.
@@ 4.1 Attention-based Operation
Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively hassanin2024visual . In diffusion models, these operations generally involve manipulating the way a model focuses on different parts of data, often through a method known as the Query-Key-Value (QKV) scheme. While large-scale pre-training has equipped this module with strong feature extraction capabilities, there remains significant ongoing work to enhance its performance for customization tasks.
Explicit attention weight manipulation. A cluster of studies focuses on restricting the influence of the SoI token within the attention layers. For example, Mix-of-Show 46gu2024mix designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity. This avoids the misalignment between the words and visual regions. DreamTuner 86hua2023dreamtuner designs a self-subject-attention layer to further refine the subject identity. This attention module takes the features of the generated image as query, the concatenation of generated features as key, and reference features as value. Layout-Control 152chen2024training adjusts attention weights specifically around the layout without additional training. Cones 2 62liu2023cones also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation. VICO 20hao2023vico inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
Implicit attention guidance. In addition to these explicit attention weights modification methods, many researchers 54xiao2023fastcomposer ; 8han2023svdiff ; 74avrahami2023break ; 16ma2023subject ; 81safaee2023clic ; 44rahman2024visual ; 90hyung2023magicapture employ localization supervision in the cross-attention module. Specifically, they train cross-attention modules using coordinate-aware loss functions, forcing attention maps to align with annotated subject positions. DreamTuner 86hua2023dreamtuner further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
@@ 4.2 Mask-guided Generation
Pixel-level mask. Benefitting from advanced segmentation methods like SAM kirillov2023segment , the SoI can be precisely isolated from the background. Based on this strategy, plenty of studies 21zhang2023personalize ; 48wang2023hifi ; 74avrahami2023break ; 81safaee2023clic ; 51lu2024object ; 95achlioptas2023stellar ; 90hyung2023magicapture ; 92tang2023retrieving ; 34li2023generate choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances. Another technique 85cai2023decoupled further adds the masked background reconstruction for better disentanglement. In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal 54xiao2023fastcomposer ; 8han2023svdiff ; 74avrahami2023break ; 16ma2023subject ; 81safaee2023clic ; 44rahman2024visual ; 90hyung2023magicapture to adjust the attention’s concentration adaptively. Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns 19wei2023elite ; 24jia2023taming ; 16ma2023subject .
Feature-level mask. In addition to the pixel-level manipulation, masks can be extended to feature-level operations. DisenBooth 18chen2023disenbooth defines an identity-irrelevant embedding with a learnable mask. By maximizing the cosine similarity between the identity-preservation embedding and identity-irrelevant embedding, the mask will adaptively exclude the redundant information, and thus the subject appearance can be better preserved. AnyDoor 14chen2023anydoor defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process. Face-Diffuser 52wang2023high determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model. Each model makes its own noise prediction, and the final noise output is a composite created through mask-guided concatenation.
@@ 4.3 Data Augmentation
Compositional augmentation. Some methods enhance data diversity through classical image augmentation like blending and spatial rearrangement. SVDiff 8han2023svdiff manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios. Such concept composition is also used in other works 16ma2023subject ; 138chen2023videodreamer ; 145wang2024customvideo . BLIP-Diffusion 10li2024blip segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset. Besides, StyleAdapter 26wang2023styleadapter chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style. PACGen 34li2023generate shows that the spatial position also entangles with the identity information. Rescaling, center crop, and relocation are effective solutions for this issue.
Synthetic data. Generated synthetic data can provide a large amount of training resources when coupled with quality assurance mechanisms. SuTI 11chen2024subject establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI. These synthetic samples then train the target PTA model. Similarly, DreamIdentity 89chen2023dreamidentity leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image. StyleDrop 5sohn2024styledrop and GAL 151zhang2024generative implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
External sources. It is intuitive to leveraging web resources to expand training datasets. COTI 3yang2023controllable adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
@@ 4.4 Regularization
Auxiliary data regularization. To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, a commonly adopted method is to use an additional dataset composed of images with the same category of the SoI 2ruiz2023dreambooth . By reconstructing these images, the personalized model is required to generate diverse instances of the class while adapting to the target subject. Building on this strategy, StyleBoost 53park2023styleboost introduces an auxiliary style-specific auxiliary data to separate content and aesthetic adaptation. Later, a dataset 57he2023data is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
Text embedding constraints. The semantic richness of pre-trained text (e.g., subject class name) provides a powerful regularization signal for personalized generation. By strategically constraining how subject-specific representations interact with established linguistic concepts in the embedding space, these approaches can achieve better generalization ability. For example, Perfusion 7tewel2023key constrains key projections toward class noun embeddings while learning value projections from subject images. Inspired by coached active learning wei2012coaching ; wei2011coached , which uses anchor concepts for optimization guidance, Compositional Inversion 9zhang2023compositional employs a set of semantically related tokens as anchors to constrain the token embedding search. In addition, some works 12gal2023encoder ; 51lu2024object regularize learnable token offsets relative to pre-trained CLIP embeddings. By minimizing the offset, the final word embedding is able to achieve better text alignment. Similarly, Cones 2 62liu2023cones minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun. And 66pang2023cross optimizes the learnable token towards the mean textual embedding of 691 well-known names. Domain-Agnostic 17arar2023domain proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples. On the other hand, VICO 20hao2023vico empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI. To leverage this discovery, an L2 loss is leveraged to reduce the difference of attention similarity logits between the SoI token and <|EOT|>.
5 Categorization of Personalization Tasks
As shown in Fig. 6, personalization covers a range of areas, including objects, styles, faces, etc. The following subsections analyze these tasks through the frameworks established in Sec. 3, examining both TTF and PTA approaches for each domain.
Figure 6: Visualization of several personalized content synthesis tasks. Given text prompts and a few images of the subject of interest, personalization approaches are required to produce expected images.

### v4

@@ 3.1.4 Prompt Engineering
The construction of training prompts for samples typically starts with adding prefix words before the modifier token. A simplest example is “Photo of V*”. However, DreamBooth [4] noted that such a simple description causes a long training time and unsatisfactory performance. To address this, they incorporate the unique modifier with a class noun to describe the SoI in the references (e.g., “Photo of V* cat”). Also, the training caption for each training reference can be more precious for better disentanglement of SoI and irrelevant concepts [40], such as “Photo of V* cat on the chair”. This follows the tending that high-quality captions in the training set could assist in further improvement of accurate text control [41].
The PTA framework has emerged as a breakthrough approach for PCS, aiming to eliminate the computational burden of per-request fine-tuning while maintaining high-quality, subject-specific generation capabilities. To achieve this goal, this approach combines large-scale pretraining with reference-aware architectures to enable single-pass personalization, as shown in Fig. 5. Based on this architecture, three critical design factors are considered to ensure practical viability: 1) Preservation of semantic-critical features to guarantee visual consistency with reference inputs, 2) Effective fusion that combines reference features with text guidance to achieve desired generation, 3) Optimization of training dataset scale to achieve robust generalization without overfitting. In the following sub-section, we first present overall architecture introduction in Section 3.1.5 and then delve into three detailed factors.
3.1.5 Pre-trianed Adaption
@@ 3.2.1 Subject Feature Extraction
Extracting representative features of the SoI is crucial in the creation of personalized content. A common approach is to employ an encoder, leveraging pre-trained models such as CLIP [42] and BLIP [43]. While these models excel at capturing global features, they often include irrelevant information that can detract from the fidelity, potentially compromising the quality of the personalized output, such as including the same background in the generation. To mitigate this issue, some studies incorporate additional prior knowledge to guide the learning process so as to focus on the targeted SoI. For instance, the SoI-specific mask [44, 32, 38, 45, 34, 35] contributes to the effective exclusion of the influence of the background. Moreover, using facial landmarks [6] in the context of human face customization helps improve identity preservation. We will discuss a more detailed technical summary of mask-assistant generation in Section 4.2.
Handling multiple input references presents another challenge but is essential for real-world deployment. This necessitates an ensemble of features from the multiple reference images to augment the framework’s adaptability. Yet, the majority of current PTA systems are limited to supporting one reference input. Some research works [6, 46] propose to average or stack features extracted from multiple references to form a composite SoI representation.
@@ 3.2.2 Subject Feature Fusion
Concatenation-based Fusion [47, 46, 5, 48, 49, 50]. This method makes the unique modifier a placeholder token to encapsulate visual subject characteristics. The placeholder token embedding, initialized with image features from visual encoders, is concatenated with text embeddings from the language model. This combined representation subsequently guides generation through standard cross-attention layers in the diffusion process, enabling basic subject-text alignment while maintaining architectural simplicity.
Cross-attention Fusion [51, 52, 6, 53, 54, 55, 56]. This paradigm extends the U-Net architecture with specialized attention mechanisms that jointly process visual and textual conditions. For example, IP-Adapter [51] introduces decoupled cross-attention layers that maintain separate query projections for image and text features. In this case, the unique modifier directly displays the plain text.
Multimodal Encoder Fusion [57, 58, 59, 60, 61]. This approach leverages powerful multimodal encoder architectures (e.g., BLIP-2 [62]) to jointly embed visual and textual subject descriptors. BLIP-Diffusion [57] exemplifies this strategy by learning a compact subject prompt embedding that fuses image patches with textual names through Q-Former modules.
Hybrid Fusion [63, 64]. Moreover, some systems integrate multiple fusion strategies. For example, Subject-Diffusion [63] integrate both concatenation and cross-attention fusion, taking advantage of the strengths of each approach to enhance the overall personalization capability.
@@ 3.2.3 Training Data
Triplet Data (Reference Image, Target Image, Target Caption). This dataset format is directly aligned with the PCS objectives, establishing a clear relation between the reference and the personalized content. However, such large-scale triplet samples are not widely available. Several strategies have been proposed to mitigate this issue: 1) Data Augmentation. Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data [57]. 2) Synthetic Sample Generation. Methods like SuTI [52] utilize multiple TTF models to generate synthetic samples, which are then paired with original references. 3) Utilizing Recognizable SoIs. Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization [65].
Dual Data (Reference Image, Reference Caption). This dataset is essentially a simplified version of the triplet format, where the personalized content is the original image itself. Such datasets are more accessible, including collections like LAION [66] and LAION-FACE [67]. However, a notable drawback is that training tends to focus more on reconstructing the reference image rather than incorporating the text prompts. Consequently, models trained on this type of data might struggle with complex prompts that require substantial modifications or interactions with objects.
@@ 3.3 Hybrid Framework
Recently, some works have started to explore the combination of TTF and PTA methods. HyperDreamBooth [26] states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details. They first develop a PTA network, followed by a subject-driven fine-tuning. Similarly, DreamTuner [39] pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction. Then an additional fine-tuning stage is conducted for fine identity preservation. In contrast, SuTI [52] first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
@@ 4 Techniques in Personalized Content Synthesis
Building upon the architectural framework discussed in Section 3, this section analyzes learning optimization techniques applicable to both frameworks. We focus on four categories: attention mechanisms, mask-guided generation, data augmentation, and regularization strategies. These methods aim to address key challenges in PCS, like enhancing subject fidelity, minimizing the interference of redundant semantics, enhancing generalization, and avoiding overfitting.
@@ 4.1 Attention-based Operation
Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively [68]. In diffusion models, these operations generally involve manipulating the way a model focuses on different parts of data, often through a method known as the Query-Key-Value (QKV) scheme. While large-scale pre-training has equipped this module with strong feature extraction capabilities, there remains significant ongoing work to enhance its performance for customization tasks.
Explicit attention weight manipulation. A cluster of studies focuses on restricting the influence of the SoI token within the attention layers. For example, Mix-of-Show [25] designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity. This avoids the misalignment between the words and visual regions. DreamTuner [39] designs a self-subject-attention layer to further refine the subject identity. This attention module takes the features of the generated image as query, the concatenation of generated features as key, and reference features as value. Layout-Control [69] adjusts attention weights specifically around the layout without additional training. Cones 2 [70] also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation. VICO [71] inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
Implicit attention guidance. In addition to these explicit attention weights modification methods, many researchers [5, 22, 38, 63, 45, 72, 37] employ localization supervision in the cross-attention module. Specifically, they train cross-attention modules using coordinate-aware loss functions, forcing attention maps to align with annotated subject positions. DreamTuner [39] further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
@@ 4.2 Mask-guided Generation
Pixel-level mask. Benefitting from advanced segmentation methods like SAM [73], the SoI can be precisely isolated from the background. Based on this strategy, plenty of studies [44, 32, 38, 45, 34, 35, 37, 74, 75] choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances. Another technique [49] further adds the masked background reconstruction for better disentanglement. In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal [5, 22, 38, 63, 45, 72, 37] to adjust the attention’s concentration adaptively. Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns [47, 76, 63].
Feature-level mask. In addition to the pixel-level manipulation, masks can be extended to feature-level operations. DisenBooth [77] defines an identity-irrelevant embedding with a learnable mask. By maximizing the cosine similarity between the identity-preservation embedding and identity-irrelevant embedding, the mask will adaptively exclude the redundant information, and thus the subject appearance can be better preserved. AnyDoor [78] defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process. Face-Diffuser [79] determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model. Each model makes its own noise prediction, and the final noise output is a composite created through mask-guided concatenation.
@@ 4.3 Data Augmentation
Compositional augmentation. Some methods enhance data diversity through classical image augmentation like blending and spatial rearrangement. SVDiff [22] manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios. Such concept composition is also used in other works [63, 80, 81]. BLIP-Diffusion [57] segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset. StyleAdapter [82] chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style. PACGen [75] shows that the spatial position entangles with the identity information. Thus rescaling, center crop, and relocation are effective augmentation solutions.
Synthetic data. Generated synthetic data can provide a large amount of training resources when coupled with quality assurance mechanisms. SuTI [52] establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI. These synthetic samples then train the target PTA model. Similarly, DreamIdentity [48] leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image. StyleDrop [24] and GAL [19] implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
External sources. It is intuitive to leveraging web resources to expand training datasets. COTI [83] adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
@@ 4.4 Regularization
Auxiliary data regularization. To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, studies start to use an additional dataset composed of images with the same category of the SoI [4]. By reconstructing these images, the personalized model is required to generate diverse instances of the class while adapting to the target subject. Building on this strategy, StyleBoost [84] introduces an auxiliary style-specific data to separate content and aesthetic adaptation. Later, a dataset [40] is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
Text embedding constraints. The semantic richness of pre-trained text (e.g., subject class name) provides a powerful regularization signal for personalized generation. By strategically constraining how subject-specific representations interact with established linguistic concepts in the embedding space, these approaches can achieve better generalization ability. For example, Perfusion [21] constrains key projections toward class noun embeddings while learning value projections from subject images. Inspired by coached active learning [85, 86], which uses anchor concepts for optimization guidance, Compositional Inversion [87] employs a set of semantically related tokens as anchors to constrain the token embedding search. In addition, some works [88, 34] regularize learnable token offsets relative to pre-trained CLIP embeddings. By minimizing the offset, the final word embedding is able to achieve better text alignment. Similarly, Cones 2 [70] minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun. And [33] optimizes the learnable token towards the mean textual embedding of 691 well-known names. Domain-Agnostic [89] proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples. On the other hand, VICO [71] empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI. To leverage this discovery, an L2 loss is leveraged to reduce the difference of attention similarity logits between the SoI token and <|EOT|>.
5 Categorization of Image Personalization Tasks
As shown in Fig. 6, personalization covers a range of areas, including objects, styles, faces, etc. The following subsections analyze these tasks through the frameworks established in Section 3, examining both TTF and PTA approaches for each domain. We also summarize these studies in Table 2, Table 3, and Table 4 to provide a clear overview and facilitate quick comparison.
Figure 6: Visualization of several personalized image synthesis tasks. Given text prompts and a few images of the subject of interest, personalization approaches are required to produce expected images.